if 20.0 grams of sodium (na) is reacted with 10.0 grams of water (h20) Wich would be the limiting reactant according to this equation. 2na+2h2o--- 2naoh+h2
Answer #1

Moles of Na = mass / molar mass = 20 / 23 = 0.870 moles

Moles of H2O = mass / molar mass = 10 / 18 = 0.56 moles

From the equation, moles of needed = moles of H2O. Since H2O is in lower quantity, it is the limiting reagent